### v1

@@
Abstract
Light penetration depth in biological tissue is limited by tissue scattering. There is an urgent need for scattering compensation in vivo focusing and imaging, particularly challenging in photon-starved scenarios, without access to the transmission side of the scattering tissue. Here, we introduce a two-photon microscopy system with Fourier-domain open-channel coupling for scattering correction (2P-FOCUS). 2P-FOCUS corrects scattering by utilizing the non-linearity of multiple-beam interference and two-photon excitation, eliminating the need for a guide star, iterative optimization, or measuring transmission or reflection matrices. We demonstrate that 2P-FOCUS significantly enhances two-photon fluorescence signals by several tens of folds when focusing through a bone sample, compared to cases without scattering compensation at equivalent excitation power. We also show that 2P-FOCUS can correct scattering over large volumes by imaging neurons and cerebral blood vessels within a 230$\times$230$\times$500 $\mu$m3 volume in the mouse brain in vitro. 2P-FOCUS could serve as a powerful tool for deep tissue imaging in bulky organisms or live animals.
keywords:
scattering correction, open channel, two-photon microscopy, deep tissue imaging
@@ 1 Introduction
Noninvasive focusing of light and imaging of objects embedded in scattering tissue are crucial for both basic biological research and clinical applications. Tissue scattering is the primary limitation for deep tissue focusing and imaging. Advanced microscopy techniques, such as multiphoton microscopy, enable noninvasive deep tissue imaging in vivo. However, as imaging depth increases, the number of ballistic photons available for multiphoton excitation at a target location decreases exponentially. Eventually, the background intensity becomes comparable to the signal in non-sparsely labeled samples[1].
@@
To increase fluorescence signal intensity in two-photon microscopy, adaptive optics has been employed to correct low-order aberrations [2, 3]. Recent cutting-edge research [4, 5] demonstrates that combining adaptive optics with multiphoton microscopy can enhance the fluorescence signal by several folds. A spatial light modulator (SLM) or a deformable mirror, positioned at the conjugate plane of the objective lens’s back aperture, is used to modulate the phase of light in the Fourier domain. However, the patterning speed of SLM is relatively slow (¡1 kHz), limiting its application to correct dynamic scattering.
Numerous techniques have been developed to correct highly spatially varying scattering [6, 7]. However, most of these techniques are designed for transmission mode, which limits their applicability in in vivo imaging within scattering tissues. In reflection mode, where the incident and detection paths are on the same side of the scattering medium, the challenge is greater. This is because both the incident light and the emitted light (back-scattered light or emitted fluorescence) are scattered by the tissue. The scattering potentials for the incident and emission light usually differ [8], necessitating distinct corrections. Secondly, in biological tissue, back-scattered light or emitted fluorescence is typically much dimmer than forward-scattered incident light, which results in a low signal-to-noise ratio in raw measurements. Third, Correcting scattering for imaging objects embedded within tissue [9] is more challenging than correcting scattering for a highly scattering but thin layer with free space between the imaging objects and the scattering media [10, 11, 12]. Without free space between the target object and the scattering media, the range of the memory effect (also called the isoplanatic patch), in which scattering properties can be considered spatially invariant, is much smaller. Consequently, compensating for scattering in reflection mode for focusing and imaging fluorescence objects inside a scattering medium presents significant challenges.
Several techniques for active focusing and imaging through scattering media with microscopy systems in reflection mode have been demonstrated [13, 14, 15, 16, 17, 18, 19]. Wavefront shaping combined with a two-photon guide star involves measuring scattered speckles using the guide star and compensating for scattering by modulating the excitation light’s phase with a deformable mirror or an SLM [13, 14, 15]. The corrective phase mask can be calculated through measurements of the electric point-spread-function [14], interference field [15], or by optimizing the two-photon signal with iterative algorithms [13]. In the absence of a guide star, techniques such as measuring the reflection matrix or fluorescence transmission matrix have been used for scattering correction in reflection mode [16, 17, 18, 20, 19]. These methods involve capturing the full input-output responses of back-scattered light via wide-field interferometric imaging and then computing a correction phase mask from the reflection matrix to counteract scattering. This approach has been integrated with two-photon microscopy for in vivo mouse brain imaging [17]. Techniques that record fluorescence-based transmission matrices or speckle patterns in reflection mode to calculate corresponding correction phase masks have also been demonstrated [18, 19]. However, most of these techniques, especially when correcting complex scattering over a large field-of-view (larger than the range of memory effect), require a significant amount of time due to the time-consuming nature of speckle pattern measurement with a camera and the computational complexity of calculating correction masks from large matrices. Furthermore, cameras can only detect bright, high-contrast speckles, which limits their application in scenarios involving dim and densely packed fluorescence objects.
To overcome current limitations, we have developed a new two-photon microscopy system named 2P-FOCUS, which stands for Fourier-domain Open-channel Coupling for Scattering correction. 2P-FOCUS modulates excitation light in the Fourier domain to couple it into high-transmission channels (‘open-channels’) within scattering tissue. Unlike previous research that identified open channels from the coherent/incoherent transmission/reflection matrix [21, 22, 23, 24, 25, 26, 27], 2P-FOCUS does so without needing to know the transmission matrix. 2P-FOCUS measures the sum of two-photon excited fluorescence as the feedback signal using a photomultiplier tube (PMT), rather than using a camera to capture scattered speckle patterns. This approach greatly reduces data collection time and computational complexity, as computing peak intensity indices is faster than calculating eigenvalues of a large transmission matrix. Since the fluorescence intensity is quadratically proportional to the excitation focus, the sum of fluorescence intensity is dominated by the intensity of the brightest speckle grain. Compared to previous open-channel research using wavefront shaping and feedback strategies [28], 2P-FOCUS uses a digital micromirror device (DMD) to modulate excitation light intensity on the relayed plane of the back aperture, rather than a SLM to modulate the phase of light. The on/off status of each DMD pixel independently controls a unique illumination angle in the real domain. The interference of multiple beams from the turned-on DMD pixels at the focus point means the enhancement of excitation intensity is quadratically proportional to the number of beams. To correct scattering outside of the range of memory effect,the DMD’s intensity modulation is synchronized with scanning mirrors to apply different correction masks to various sub-regions during imaging, which increases the system’s degree of freedom. The downside of intensity modulation is that laser power on the DMD’s off pixels, corresponding to the closed channels in the scattering media, is blocked. However, given that the power of many commercially available lasers far exceeds the thermal damage threshold for biological tissues, laser power efficiency is not a major concern. In addition, 2P-FOCUS differs from other scattering correction systems using a DMD in the Fourier domain for wavefront shaping [29, 30, 31, 32, 33, 34] in that it is a two-photon microscopy system in reflection mode for imaging embedded objects, rather than a one-photon system in transmission mode. 2P-FOCUS also differs from two-photon microscopy systems that use an intensity modulation in the real domain for scattering correction via post-processing [35, 36, 37, 38, 39, 40], as 2P-FOCUS actively corrects scattering to deliver more laser energy through scattering media. By leveraging the non-linearity of multiple beam interference and two-photon excitation, 2P-FOCUS results in fluorescence intensity enhancement being proportional to the fourth power of the number of interference beams.
We present the experimental setup and computational methods of 2P-FOCUS for focusing and imaging tasks with various biological samples. We demonstrate that 2P-FOCUS achieves approximately a 63-fold enhancement in fluorescence intensity compared to uncorrected scenarios. Additionally, 2P-FOCUS can correct scattering over a large field-of-view by projecting different correction masks onto various sub-regions, enabling correction beyond the memory effect range. We demonstrate two-photon imaging of fluorescence-labeled neurons and cerebral blood vessels up to 500 $\mu$m deep over a $230\times 230\mu m^{2}$ field-of-view in the mouse brain in vitro. 2P-FOCUS effectively and rapidly corrects scattering, potentially making it a powerful tool for intravital two-photon imaging in vivo.
2 Results
Figure 1: Principle and schematic diagram of two-photon microscopy with Fourier-domain open-channel coupling for scattering correction (2P-FOCUS). (a) The case Without correction. All pixels of the digital micromirror device (DMD) are turned on to pass all incident beams to the objective lens and the scattering sample. These beams are scattered and cannot form a tight focus. (b) The process of generating a correction mask. The first step is to project random intensity patterns, which are orthogonal to each other, onto the DMD and record the corresponding fluorescence intensity using a photomultiplier tube (PMT). The second step is to select the random masks that correspond to high fluorescence intensity and summing these masks to create a grayscale correction mask. The third step is the binarization of the grayscale correction mask for DMD display. (c) The case after correction. By projecting the binary correction mask, only the beams that interfere constructively and generate a bright focus are selected. The illumination power on the sample is maintained at the same level before and after correction by increasing the input power to the DMD to compensate for the power loss due to turning off some pixels. The correction results in a brighter focus compared to the case before correction. (d) Optical schematic diagram of 2P-FOCUS. A femtosecond laser at 1035 nm is used as the light source. A half-wave plate (HWP) and a polarized beam splitter (PBS) are used to control the laser power. A pair of scanning mirrors (SM) is used to scan the focus across the field-of-view. A grating is used to compensate for the spatial dispersion caused by the DMD. L1-L2, L3-L4, and L5-tube lens (TL) are 4-f relay lenses. DMD is located on the conjugate plane of the objective lens’s (OL) back aperture. The objective lens focuses the beams on the sample and collects emitted fluorescence. Fluorescence is detected by a PMT after passing through a dichroic mirror (DM), a collection lens, and a color filter. Sample is mounted on a linear translation stage for collecting z-stack images.

### v2

@@
Scattering Correction through Fourier-Domain Intensity Coupling in Two-Photon Microscopy (2P-FOCUS)
[…]
journal: opticajournalarticletype: Research Article
{abstract*}
Light penetration depth in biological tissue is limited by tissue scattering. Correcting scattering becomes particularly challenging in scenarios with limited photon availability and when access to the transmission side of the scattering tissue is not possible. Here, we introduce a new two-photon microscopy system with Fourier-domain intensity coupling for scattering correction (2P-FOCUS). 2P-FOCUS corrects scattering by intensity modulation in the Fourier domain, leveraging the nonlinearity of multiple-beam interference and two-photon excitation, eliminating the need for a guide star, iterative optimization, or measuring transmission or reflection matrices. 2P-FOCUS uses random patterns to probe scattering properties, combined with a single-shot algorithm to rapidly generate the correction mask. 2P-FOCUS can also correct scattering beyond the limitation of the memory effect by automatically customizing correction masks for each subregion in a large field-of-view. We provide several proof-of-principle demonstrations here, including focusing and imaging through a bone sample, and imaging neurons and cerebral blood vessels in the mouse brain ex vivo. 2P-FOCUS significantly enhances two-photon fluorescence signals by several tens of folds compared to cases without scattering correction at the same excitation power. 2P-FOCUS can also correct tissue scattering over a 230$\times$230$\times$510 $\mu$m3 volume, which is beyond the memory effect range. 2P-FOCUS is able to measure, calculate, and correct scattering within a few seconds, effectively delivering more light deep into the scattering tissue. 2P-FOCUS could be broadly adopted for deep tissue imaging owing to its powerful combination of effectiveness, speed, and cost.
@@ 1 Introduction
Noninvasive focusing of light and imaging of objects embedded in scattering tissue are crucial for both basic biological research and clinical applications. Tissue scattering is the primary limitation for deep tissue focusing and imaging. Advanced microscopy techniques, such as confocal microscopy and multiphoton microscopy, enable noninvasive deep tissue imaging. However, as imaging depth increases, the number of ballistic photons available for excitation at a target location decreases exponentially. Eventually, the background intensity becomes comparable to the signal in non-sparsely labeled samples [1].
@@
To increase fluorescence signal intensity in two-photon microscopy, adaptive optics has been employed to correct low-order aberrations [2, 3]. Recent cutting-edge research [4, 5] demonstrates that combining adaptive optics with multiphoton microscopy can enhance the fluorescence signal by several folds. A spatial light modulator (SLM) or a deformable mirror, positioned at the conjugate plane of the objective lens’s back aperture, is used to modulate the phase of light in the Fourier domain. However, the patterning speed of SLM is relatively slow ($<$1 kHz), limiting its application to correct dynamic scattering [6].
Numerous techniques have been developed to correct highly spatially varying scattering [7, 8, 9, 10, 11]. However, most of these techniques are designed for transmission mode, which limits their applications for imaging objects embedded within the scattering tissues of living animals. Among the reflection mode techniques, many are post-processing methods [12, 13, 14, 15, 16], including our previous work [17, 18, 19, 20, 21], which require significant time for image reconstruction ranging from several minutes to over an hour. Unlike post-processing techniques that correct scattering in images offline, several techniques achieve active correction of scattering by wavefront shaping using an SLM in reflection-mode multiphoton microscopy systems [22, 23, 24, 25]. The wavefront shaping is based on interferometry measurements. Excitation light is split into two beams, either in a common path [22, 24] or in a divergent path [23, 25]. One beam is steady, while the other beam is phase-shifted using either an SLM or a galvo-scanner. The interference of these two beams changes with the phase shifts, resulting in changes in fluorescence intensity, which is used as the feedback signal. The phase mask can then be calculated and optimized after a few iterations. Even though these techniques achieve deep tissue imaging, they still face some limitations. First, the maximum improvement in fluorescence intensity for brain imaging is usually less than tenfold [23, 24, 25]. Second, the field-of-view after scattering correction is usually limited to a few tens of microns in brain imaging due to the limited range of the memory effect (i.e., isoplanatic patch, mean free path). Third, the total correction speed is relatively slow, including not only the time for taking measurements but also the time for computing the mask, projecting the mask, and reloading updated phase patterns to the SLM for iterative optimization. A detailed comparison between these methods and our technique is summarized in Appendix Table S1.
Active correction of scattering via intensity modulation using a digital micromirror device (DMD) enables high-speed scattering correction. “Intensity modulation” here refers to techniques that directly modulate the intensity of light at the Fourier plane as opposed to “phase modulation” discussed above which directly modulates the phase of light at the Fourier plane. Most commercially available DMDs have patterning speeds ranging from tens to over 30 kHz, which is significantly faster than the state-of-the-art SLMs. DMDs are also generally more cost-effective than SLMs. Intensity modulation for scattering correction has been demonstrated by focusing a laser beam through a scattering medium and detecting the transmitted laser light with a detector [26, 27, 28, 29, 30, 31, 32]. The correction mask can be probed by turning on segments on the intensity modulator one-by-one [26], by projecting Hadamard patterns to calculate transmission matrix [27, 28], or by projecting random patterns combined with transmission matrix measurement [29] or iterative optmization algorithms [30, 31, 32]. Compared to phase modulation, traditional intensity modulation faces a challenge: for a given incident laser power on the modulator, scattering correction increases the signal, but intensity modulation reduces the signal because the total intensity on the detector decreases with the turned-off pixels. This is not an issue when focusing a laser beam through scattering media and detect the transmitted laser light [26, 27, 28, 29, 30, 31, 32], as the photon flux from a laser is high.
However, in fluorescence imaging, especially multiphoton imaging of biological samples, scattering correction by intensity modulation is very challenging because: (1) The number of fluorescent photons under two-photon excitation is much lower than that of a laser beam, and the maximum excitation intensity is constrained by the thermal damage threshold of biological samples; (2) Fluorescent photons are collected in reflection mode rather than in transmission mode, while most light is forward-scattered by biological tissues, further reducing the number of photons after scattering; (3) Two-photon microscopy images embedded objects within scattering tissue, without any free space between the phantom and the focus, which limits the degrees of freedom for correcting scattering [33, 34]; (4) Imaging a large field of view requires scattering correction beyond the memory effect range, which is more difficult than correcting scattering at a single spot.
Recently, high-power femtosecond lasers (ranging from a few watts to over 100 W) for multiphoton microscopy have become commercially available, making two-photon microscopy with intensity modulation feasible [35, 36]. It has been used for aberration correction [35, 36], but not yet for scattering correction because the challenges mentioned above remain unsolved. In these work [35, 36] for aberration correction, the DMD is used as a wavefront modulator by calculating binary holographic masks using Lee holography [37], with Zernike modes employed as the orthogonal basis to probe aberrations. The correction mask is generated using a modified hill-climbing algorithm similar to the classic algorithm in adaptive optics [38], where the final correction mask is a linear combination of weighted Zernike modes, and the weight coefficient of each Zernike mode is iteratively measured. However, binary intensity modulation generated by Lee holography is less efficient and effective than phase modulation, which has been demonstrated for focusing through scattering media [39, 29]. Consequently, while this technique corrects aberrations in two-photon microscopy, fluorescence intensity is only improved by 2.5 to 4 times when imaging ex vivo Drosophila brains at depths of 100 - 200 $\mu$m. Despite the DMD’s fast operational speed of 22.7 kHz, the iterative hill-climbing algorithm is slow, leading to a total processing time of approximately 6 minutes. Therefore, instead of adapting orthogonal bases (i.e., Zernike modes) and algorithms (i.e., iterative hill-climbing algorithm) designed for phase modulation, innovative approaches tailored specifically for two-photon microscopy with binary intensity modulation are needed, including new orthogonal bases and algorithms. While the key advantage of DMD-based intensity modulation is its patterning speed, no two-photon microscopy system with binary intensity modulation has yet been able to rapidly and effectively correct scattering in deep tissue.
To overcome these challenges, we have developed a new two-photon microscopy system named 2P-FOCUS (FOurier-domain intensity Coupling for Scattering correction) to provide effective and fast scattering correction over a large field of view. The power loss from intensity modulation is compensated by increasing the laser power input to the intensity modulator (i.e., a DMD) to maintain the same laser power on the sample before and after modulation. Unlike phase modulation, intensity modulation segments the incident collimated beam into multiple narrow beams, generating multiple-beam interference within the scattering tissue. To select the beams that remain in phase after scattering while blocking those that are out of phase, we use random patterns to probe whether the randomly selected beams are in phase or not after scattering. In-phase beams generate a bright main lobe after interference, resulting in brighter fluorescence compared to non-constructive interference. The final correction mask is generated by summing those random patterns followed by binarization. Random patterns are used instead of patterns at specific frequencies (e.g., Hadamard [27, 28] or Zernike patterns [35, 36]) because: (1) Each random pattern consists of a broad range of spatial frequencies, allowing us to multiplex measurements across different frequencies; (2) random patterns can flexibly control the ratio of turned-on pixels to the total number of pixels (“sparsity") to improve the correction but Hadamard or Zernike patterns cannot; and (3) random patterns generate a central focus with stable intensity and location, whereas Hadamard patterns produce multiple foci that vary in intensity and position with each pattern. This single-shot algorithm ensures constructive interference without the need for time-consuming iterative optimization. To correct scattering beyond the memory effect range, we synchronize the DMD with a galvo-scanner on its conjugate plane to project different correction masks onto corresponding regions during scanning. This strategy significantly increases the system’s degrees of freedom and effectively controlling a 4D light field [40].
We demonstrate the impressive capabilities of 2P-FOCUS through a series of proof-of-principle experiments. These include focusing and imaging through approximately 200 $\mu$m-thick bone and capturing fluorescence-labeled neurons and blood vessels at depths of up to 510 $\mu$m deep in ex vivo mouse brain tissue. Remarkably, 2P-FOCUS achieves over a 36-fold increase in fluorescence intensity compared to standard two-photon microscopy when imaging in the brain tissue. 2P-FOCUS also achieves a 230 $\times$ 230 $\mu m^{2}$ field-of-view in all experiments—beyond the memory effect range and significantly larger than that of current methods. Furthermore, 2P-FOCUS completes the entire correction process in just a few seconds, including the time for measurements, calculation, and projection of correction masks, far surpassing the speed of similar techniques [35, 36]. To our knowledge, 2P-FOCUS is the first two-photon microscopy system to use intensity modulation for rapid and effective scattering correction.
2 Methods